OpenStudy (anonymous):

Choose the correct first step(s) to solve the system using elimination. 4x + 7y = 6 and 6x + y = 2 7(6x + y = 2) → -42x - 7y = -14 4x + 7y = 6x + y y = -6x + 20 → 4x + 7(-6x + 20) = 6 (4x + 7y) + (6x + y) = 8 (4x + 7y) - (6x + y) = 4

OpenStudy (anonymous):

in terms of x—1st & 2nd equations: 4x + 7y + 6x + y = 6 + 2 10x + 8y = 8 y = 1 - 5/4x 1st equation—x, substitute y: 4x + 7(1 - 5/4x) = 6 4x + 7 - 35/4x = 6 16x + 28 - 35x = 24 19x = 4 x = 4/19 Value of y, substitute x: = 1 - 5/4(4/19) = 19/19 - 5/19 = 14/19 Answer: x = 4/19, y = 14/19 Proof—2nd equation: 6(4/19) + 14/19 = 2 24/19 + 14/19 = 2 38/19 = 2 2 = 2

OpenStudy (anonymous):

i think there may be a typo here

OpenStudy (anonymous):

if the initial equations were \[4x + 7y = 6 \] \[6x + y = 20\] the a good first step would be \[y = -6x + 20 → 4x + 7(-6x + 20) = 6 \]